OpenStudy (anonymous):

Explain how the formula for the area of a triangle is derived from that of a rectangle.

OpenStudy (anonymous):

Look at the formulas: Triangles : A = 1/2 * B * H Rectangles : A = L * W

OpenStudy (anonymous):

It is basically the same thing, except that triangles are halved.

OpenStudy (unklerhaukus):